# Analytic theory of dropout regularization

### v1

@@ Abstract
Dropout is a regularization technique widely used in training artificial neural networks to mitigate overfitting. It consists of dynamically deactivating subsets of the network during training to promote more robust representations. Despite its widespread adoption, dropout probabilities are often selected heuristically, and theoretical explanations of its success remain sparse. Here, we analytically study dropout in two-layer neural networks trained with online stochastic gradient descent. In the high-dimensional limit, we derive a set of ordinary differential equations that fully characterize the evolution of the network during training and capture the effects of dropout. We obtain a number of exact results describing the generalization error and the optimal dropout probability at short, intermediate, and long training times. Our analysis shows that dropout reduces detrimental correlations between hidden nodes, mitigates the impact of label noise, and that the optimal dropout probability increases with the level of noise in the data. Our results are validated by extensive numerical simulations.
@@ 1 Introduction
More than a decade after its introduction, dropout remains a widely used technique in training state-of-the-art models [3], with several generalizations proposed. One such variant is DropConnect [4], which applies dropout at the level of individual weights rather than entire nodes. While early studies [1, 2] suggested an optimal activation probability between 0.5 and 0.8, constant over training, more recent work has advocated for adaptive schedules [5, 6].
@@
In contrast to these prior approaches, in this paper we focus on the typical-case scenario and derive an exact analytic characterization of online feature-learning dynamics in the presence of dropout regularization. To this end, we leverage tools from statistical physics to obtain closed-form equations for the dropout dynamics in two-layer networks [16, 17, 18, 19] within a prototypical teacher-student framework [20, 21, 22]. This theoretical framework allows us to address several open questions about the mechanisms driving the performance improvement induced by dropout. In particular, we clarify the benefits of dropout in the presence of dataset noise, identify optimal dropout probabilities, and analyze its impact on the different phases of training.
@@ 2.1 Model definition
$\phi^{*}(x)=\sum_{m=1}^{M}v^{*}_{m}\,g^{*}\left(\frac{w^{*}_{m}\cdot x}{\sqrt{N}}\right)\,,$ (2)
where $v^{*}\in\mathbb{R}^{M}$, $w^{*}_{m}\in\mathbb{R}^{N}$, and $g^{*}$ denote the fixed teacher weights and activation function. Here the symbol $\cdot$ denotes the scalar product. To learn the dataset, we employ a student network, which after $\mu$ training steps takes the form
$\phi^{\mu}(x)=\sum_{i=1}^{K}\sigma^{\mu}_{i}v^{\mu}_{i}g\left(\frac{w^{\mu}_{i}\cdot x}{\sqrt{N}}\right)\,,$ (3)
@@
We assume that these variables are independently drawn from a Bernoulli distribution with activation probability $0<r\leq 1$, i.e., $\sigma_{i}^{\mu}\sim\operatorname{Bernoulli}(r)$. These binary variables model the effect of dropout: at each training step $\mu$, each hidden node is active with probability $r$ or dropped out with probability $1-r$.
@@
where $\eta$ and $\eta_{v}$ are the learning rates for the first and second layers, respectively. Here, the updates are evaluated on an independent data point $\{x^{\mu},y^{\mu}\}$ freshly drawn at each step. Note that this setup differs from offline (or batch) learning, where training examples are typically reused multiple times. We have defined the error term $\Delta^{\mu}=\hat{y}^{\mu}-y^{\mu}$ and the pre-activations $\lambda_{k}^{\mu}=w_{k}^{\mu}\cdot x^{\mu}/\sqrt{N}$. This type of dynamics was previously investigated in Refs. [16, 17, 19], where closed equations for low-dimensional order parameters were derived. Here, we would like to extend these results to the case of dropout.
As customary when using dropout, at testing time the full network $\phi_{f}(x)$ is used, after a rescaling of the weights by a factor $0<r_{f}\leq 1$:
$\phi_{f}(x)=\sum_{i=1}^{K}r_{f}\,v^{\mu}_{i}\,g\left(\frac{w^{\mu}_{i}\cdot x}{\sqrt{N}}\right)\,.$ (7)
@@
where we use the convention that the indices $i,j,k,\ell$ refer to the student network (running from $1$ to $K$), while $m,n$ refer to the teacher (running from $1$ to $M$). Unless stated otherwise, we assume $T_{nm}=\delta_{nm}$. The order parameters provide physical intuition into the learning dynamics. For instance, $R_{in}$—a.k.a. the magnetization—quantifies the alignment between the $i$-th student node and the $n$-th teacher node.
@@
$\begin{split}\epsilon_{g}&=\frac{r_{f}^{2}}{\pi}\sum_{i,k}v_{i}v_{k}\arcsin\left(\frac{Q_{ik}}{\sqrt{1+Q_{ii}}\sqrt{1+Q_{kk}}}\right)\\ &\quad+\frac{1}{\pi}\sum_{n,m}v_{n}^{*}v_{m}^{*}\arcsin\left(\frac{T_{nm}}{\sqrt{1+T_{nn}}\sqrt{1+T_{mm}}}\right)-\frac{2r_{f}}{\pi}\sum_{i,n}v_{i}v_{n}^{*}\arcsin\left(\frac{R_{in}}{\sqrt{1+Q_{ii}}\sqrt{1+T_{nn}}}\right).\end{split}$ (11)
@@
Figure 1: Comparison of learning dynamics with and without dropout. Left: Generalization error vs. training time $\alpha=\mu/N$ with ($r=0.68$) and without ($r=1$) dropout. Center: The observable $\tilde{\Delta}=(Q_{12}-R_{11}R_{21})/\sqrt{Q_{11}Q_{22}}$, which quantifies detrimental correlations between the student’s hidden nodes. Right: Cosine similarity between the student’s first hidden node weight vector $w_{1}$ and the teacher weight vector $w^{*}_{1}$. Parameters: $K=2$, $M=1$, $\eta=1$, $\eta_{v}=0$, $\alpha=5$, and $\sigma=0.3$. The teacher weights $w^{*}$ are drawn i.i.d. from $\mathcal{N}(0,1)$; the student weights are initialized i.i.d. from $\mathcal{N}(0,0.1)$, with $N=50000$. The second-layer weights are fixed to one. Solid lines show theoretical predictions from the ODEs in Eq. (10); crosses mark numerical simulations on a single training trajectory.
@@ 2.2 Summary of the main results
  • Analytic theory of dropout: We develop an analytic framework describing the online SGD learning dynamics with dropout in a prototypical teacher-student model. The dynamics are captured by a set of ODEs, given in Eq. (10), which are exact in the high-dimensional limit $N\to\infty$. While these ODEs cannot be solved analytically in general, we first examine their numerical integration and then explore specific limits where analytical progress can be made.
  • Numerical analysis of dropout: As a first step, we integrate Eq. (10) numerically to investigate the effects of dropout on the high-dimensional learning dynamics (see Fig. 1). We observe that, although dropout decreases the cosine similarity between student and teacher nodes, it leads to more independent student representations, resulting in a substantial improvement in performance. In particular, we find that dropout mitigates the negative effects of label noise. Indeed, the optimal activation probability $r^{*}$ decreases with the variance $\sigma^{2}$ of the label noise (see Fig. 2).
  • Early-time dynamics: At the beginning of training, we assume that the student weights are initialized with a small norm and expand the dynamics to leading order in that norm. The evolution equations can be solved exactly for arbitrary $M$, $K$, and $r$. As a result, we compute the optimal activation probability that minimizes the generalization error:
    $\begin{split}r^{*}=\min\left\{1;\frac{12M}{3M\eta+9\eta\sigma^{2}+\sqrt{3\eta(M+3\sigma^{2})\left[M(32(K-1)+3\eta)+9\eta\sigma^{2}\right]}}\right\}\,.\end{split}$ (12)
    This expression matches well with the numerical result shown in Fig. 2. Interestingly, $r^{*}$ decreases with increasing noise variance $\sigma^{2}$, consistent with the intuition that dropout prevents noise fitting. It also decreases with an increasing number of hidden nodes $K$, reflecting the role of dropout in mitigating negative co-adaptations between hidden units. Furthermore, we obtain the phase diagram in Fig. 3, which identifies the regimes in which dropout improves generalization as a function of the problem parameters.
  • Unspecialized phase (low-$\eta$ limit): We find that the generalization error quickly reaches a plateau, corresponding to an unspecialized phase in which each student node overlaps equally with all teacher nodes. This unspecialized phase is known to be unstable in the absence of dropout [18]. The plateau corresponds to a fixed point of the ODEs, which cannot be computed analytically in general. However, in the limit of small learning rate $\eta$, it is possible to derive an analytic expression for this fixed point. In this regime, we obtain the following exact expression for the generalization error (see also Fig. 4)
@@
    $\begin{split}R_{0}=\frac{1}{\sqrt{\left(1+(K-1)r\right)^{2}-M+\sqrt{M^{2}+(K-1)r\left(1+(K-1)r\right)^{2}\left(2+(K-1)r\right)}}}\,.\end{split}$ (14)
    The optimal activation probability $r_{0}^{*}$ is obtained by minimizing the expression for the generalization error and is shown in Fig. 5. We observe that the activation probability decreases with increasing $K$. Note, however, that in the ODEs (10), the dependence on the label noise variance $\sigma^{2}$ disappears at leading order in $\eta$. This motivates us to compute the first-order correction in $\eta$ to determine how the optimal dropout probability depends on $\sigma^{2}$.
  • Unspecialized phase (small-$\eta$ corrections): To investigate the role of label noise in determining the optimal activation probability, we derive perturbative corrections to the fixed point to leading order in $\eta$. In particular, we show analytically that the optimal activation probability takes the form
@@
    where we compute the coefficients $a(K,M)$ and $b(K,M)$ explicitly (see Fig. 6). Notably, we find that $b(K,M)<0$ for all values of $K$ and $M$ considered, confirming that the optimal activation probability decreases with increasing label noise.
  • Specialization transition: In the absence of dropout ($r=1$), the unspecialized phase is known to be linearly unstable [18], leading to specialization at long times—i.e., each student node aligns preferentially with one of the teacher nodes. Here, we investigate how dropout modifies this specialization transition. By performing a linear stability analysis around the plateau fixed point, we derive the phase diagram shown in Fig. 7, valid for $K=M$ and small $\eta$. Interestingly, we find that specialization occurs only for sufficiently large values of $r$ (i.e., low dropout). Below a critical threshold in $r$, the plateau phase becomes stable and the system does not transition to specialization.
@@ 3 Results
In this section, we begin by discussing general insights obtained from the numerical integration of the ODEs (10). We then proceed with an analytical study of these equations, deriving explicit results for the optimal activation probability and the corresponding generalization performance across the different phases of the learning dynamics.
@@ 3.1 Numerical analysis of the dropout dynamics
We can therefore identify two competing objectives: 1) learning the signal $w_{1}^{*}$ effectively (favored by a high activation probability) and 2) avoiding fitting the noise by promoting independence between $\tilde{w}_{1}$ and $\tilde{w}_{2}$ (favored by a low activation probability).
@@
Interestingly, dropout reduces the alignment of the weights with the signal $w_{1}^{*}$ (right panel in Fig. 1). This is expected as dropout effectively reduces the number of training updates for each individual node. Despite the weaker signal representation, the increased independence among nodes outweighs this drawback, leading to superior generalization.
A key parameter in our setting is the variance $\sigma^{2}$ of the label noise. When training without dropout, the network is expected to be more prone to fitting this noise, which can degrade performance. In practical applications, dropout is commonly observed to mitigate noise fitting and promote a more robust representation of the data.
Figure 2: Optimal activation probability $r^{*}$ vs. the variance $\sigma^{2}$ of the label noise. Parameters: $K=2$, $M=1$, $\eta=1$, $\eta_{v}=0$, $\alpha=0.1$. The teacher weight vector $w^{*}$ is taken to be of unit norm, the student weights $w$ are initialized to zero. The second layer weights are fixed to one. The continuous line represents the optimal probability obtained by numerical optimization by using the full equations of motion. The crosses correspond to the analytical result in Eq. (23), obtained from a short-time expansion.
In Fig. 2 we examine the optimal dropout probability as a function of the label-noise variance $\sigma^{2}$. Here, the crosses are obtained by numerical optimization of the full dynamical equations, while the continuous line corresponds to the asymptotic result derived analytically in Sec. 3.2. In the absence of label noise ($\sigma=0$), the optimal activation probability $r^{*}$ remains close to unity, indicating minimal dropout. Conversely, as the label noise increases, the optimal strategy shifts toward a lower activation probability, corresponding to a higher dropout probability.
As we show in the following sections, the fact that the optimal activation probability $r^{*}$ decreases with increasing label noise holds robustly across the different phases of the learning dynamics.

### v2

@@ Abstract
Dropout is a regularization technique widely used in training artificial neural networks to mitigate overfitting. It consists of dynamically deactivating subsets of the network during training to promote more robust representations. Despite its widespread adoption, dropout rates are often selected heuristically, and theoretical explanations of its success remain sparse. Here, we analytically study dropout in two-layer neural networks trained with online stochastic gradient descent. In the high-dimensional limit, we derive a set of ordinary differential equations that fully characterize the evolution of the network during training and capture the effects of dropout. We obtain a number of exact results describing the generalization error and the optimal dropout probability at short, intermediate, and long training times. Our analysis shows that dropout reduces detrimental correlations between hidden nodes, mitigates the impact of label noise, and that the optimal dropout probability increases with the level of noise in the data. Our results are validated by extensive numerical simulations.
@@ 1 Introduction
More than a decade after its introduction, dropout remains a widely used technique in training state-of-the-art models [3], with several generalizations proposed. One such variant is DropConnect [4], which applies dropout at the level of individual weights rather than entire nodes. While early studies [1, 2] suggested an optimal activation rate between 0.5 and 0.8, constant over training, more recent work has advocated for adaptive schedules [5, 6].
@@
In contrast to these prior approaches, in this paper we focus on the typical-case scenario and derive an exact analytic characterization of online feature-learning dynamics in the presence of dropout regularization. To this end, we leverage tools from statistical physics to obtain closed-form equations for the dropout dynamics in two-layer networks [16, 17, 18, 19] within a prototypical teacher-student framework [20, 21, 22]. This theoretical framework allows us to address several open questions about the mechanisms driving the performance improvement induced by dropout. In particular, we clarify the benefits of dropout in the presence of dataset noise, identify optimal dropout rates, and analyze its impact on the different phases of training.
@@ 2.1 Model definition
Figure 1: a) Illustration of the model presented in Section 2.1. b) Sketch of the generalization error dynamics highlighting the three phases of learning. At early times, the weights remain close to their small-norm initialization. During the plateau (unspecialized phase), the hidden units are symmetric, being equally aligned with the teacher units. After specialization, this symmetry is broken.
@@
$\phi^{*}(x)=\sum_{m=1}^{M}v^{*}_{m}\,g^{*}\left(\frac{w^{*}_{m}x}{\sqrt{N}}\right)\,$, (2)
where $v^{*}\in\mathbb{R}^{M}$, $w^{*}_{m}\in\mathbb{R}^{N}$, and $g^{*}$ denote the fixed teacher weights and activation function. To learn the dataset, we employ a student network, which after $\mu$ training steps takes the form
$\phi^{\mu}(x)=\sum_{i=1}^{K}\sigma^{\mu}_{i}v^{\mu}_{i}g\left(\frac{w^{\mu}_{i}x}{\sqrt{N}}\right)\,$, (3)
@@
We assume that these variables are independently drawn from a Bernoulli distribution with activation probability $0<r\leq 1$, i.e., $\sigma_{i}^{\mu}\sim\operatorname{Bernoulli}(r)$. These binary variables model the effect of dropout: at each training step $\mu$, each hidden node is active with probability $r$ or dropped out with probability 1r1-r. A pictorial representation of this model is shown in Figure 1a.
@@
where $\eta$ and $\eta_{v}$ are the learning rates for the first and second layers, respectively. Here, the updates are evaluated on an independent data point $\{x^{\mu},y^{\mu}\}$ freshly drawn at each step. Note that this setup differs from offline (or batch) learning, where training examples are typically reused multiple times. We have defined the error term $\Delta^{\mu}=\hat{y}^{\mu}-y^{\mu}$ and the pre-activations $\lambda_{k}^{\mu}=w_{k}x^{\mu}/\sqrt{N}$, for each time step $\mu$. This type of dynamics was previously investigated in Refs. [17, 16, 19], where closed equations for low-dimensional order parameters were derived. Here, we would like to extend these results to the case of dropout.
As customary when using dropout, at testing time the full network $\phi_{f}(x)$ is used, after a rescaling of the rates by a factor $0<r_{f}\leq 1$:
$\phi_{f}(x)=\sum_{i=1}^{K}r_{f}\,v^{\mu}_{i}\,g\left(\frac{w^{\mu}_{i}x}{\sqrt{N}}\right)\,.$ (7)
@@
where the symbol $\cdot$ denotes the scalar product and we use the convention that the indices $i,j,k,\ell$ refer to the student network (running from $1$ to $K$), while $m,n$ refer to the teacher (running from $1$ to $M$). Unless stated otherwise, we will assume $T_{nm}=\delta_{nm}$. The order parameters provide physical intuition into the learning dynamics. For instance, $R_{in}$—a.k.a. the magnetization—quantifies the alignment between the $i$-th student node and the $n$-th teacher node.
@@
$\begin{split}\epsilon_{g}&=\frac{r_{f}^{2}}{\pi}\sum_{i,k}v_{i}v_{k}\arcsin\left(\frac{Q_{ik}}{\sqrt{1+Q_{ii}}\sqrt{1+Q_{kk}}}\right)\\ &\quad+\frac{1}{\pi}\sum_{n,m}v_{n}^{*}v_{m}^{*}\arcsin\left(\frac{T_{nm}}{\sqrt{1+T_{nn}}\sqrt{1+T_{mm}}}\right)\\ &\quad-\frac{2r_{f}}{\pi}\sum_{i,n}v_{i}v_{n}^{*}\arcsin\left(\frac{R_{in}}{\sqrt{1+Q_{ii}}\sqrt{1+T_{nn}}}\right).\end{split}$ (11)
@@
Figure 2: Comparison of learning dynamics with and without dropout. Left: Generalization error vs. training time $\alpha=\mu/N$ with ($r=0.68$) and without ($r=1$) dropout. Center: The observable $\tilde{\Delta}=(Q_{12}-R_{11}R_{21})/\sqrt{Q_{11}Q_{22}}$, which quantifies detrimental correlations between the student’s hidden nodes (see main text for details). Right: Cosine similarity between the student’s first hidden node weight vector $w_{1}$ and the teacher weight vector $w^{*}_{1}$. Parameters: $K=2$, $M=1$, $\eta=1$, $\eta_{v}=0$, $\alpha=5$, and $\sigma=0.3$. The teacher weights $w^{*}$ are drawn i.i.d. from $\mathcal{N}(0,1)$; the student weights are initialized i.i.d. from $\mathcal{N}(0,0.1)$, with $N=50000$. The second-layer weights are fixed to one. Solid lines show theoretical predictions from the ODEs in Eq. (10); crosses mark numerical simulations on a single training trajectory.
@@ 2.2 Summary of the main results
  • Analytic theory of dropout: We develop an analytic framework describing the online SGD learning dynamics with dropout in a prototypical teacher-student model. The dynamics are captured by a set of ODEs, given in Eq. (10), which are exact in the high-dimensional limit $N\to\infty$. While these ODEs cannot be solved analytically in general, we first examine their numerical integration and then explore specific limits where analytical progress can be made. An illustration of the different learning phases under consideration is provided in Figure 1b.
  • Numerical analysis of dropout: As a first step, we integrate Eq. (10) numerically to investigate the effects of dropout on the high-dimensional learning dynamics (see Fig. 2). We observe that, although dropout decreases the cosine similarity between student and teacher nodes, it leads to more independent student representations, resulting in a substantial improvement in performance. In particular, we find that dropout mitigates the negative effects of label noise. Indeed, the optimal activation rate $r^{*}$ decreases with the variance $\sigma^{2}$ of the label noise (see Fig. 3).
  • Early-time dynamics: At the beginning of training, we assume that the student weights are initialized with a small norm and expand the dynamics to leading order in that norm. The evolution equations can be solved exactly for arbitrary $M$, $K$, and $r$. As a result, we compute the optimal activation rate that minimizes the generalization error:
[…]
    and matches well with the numerical result shown in Fig. 3. Interestingly, $r^{*}$ decreases with increasing noise variance $\sigma^{2}$, consistent with the intuition that dropout prevents noise fitting. It also decreases with an increasing number of hidden nodes $K$, reflecting the role of dropout in mitigating negative co-adaptations between hidden units. Furthermore, we obtain the phase diagram in Fig. 4, which identifies the regimes in which dropout improves generalization as a function of the problem parameters.
  • Unspecialized phase (analyzed in the low-$\eta$ limit): We find that the generalization error quickly reaches a plateau, corresponding to an unspecialized phase in which each student node overlaps equally with all teacher nodes. This unspecialized phase, observed numerically across different values of the learning rate, is known to be unstable in the absence of dropout [18]. The plateau corresponds to a fixed point of the ODEs, which cannot be computed analytically in general. However, in the limit of small learning rate $\eta$, it is possible to derive an analytic expression for this fixed point. In this regime, we obtain the following exact expression for the generalization error (see also Fig. 5)
@@
    $\begin{split}\frac{1}{R_{0}}=\sqrt{\left(1+(K-1)r\right)^{2}-M+\sqrt{M^{2}+(K-1)r\left(1+(K-1)r\right)^{2}\left(2+(K-1)r\right)}}\,.\end{split}$ (14)
    The optimal activation rate $r_{0}^{*}$ is obtained by minimizing the expression for the generalization error and is shown, together with the corresponding generalization error, in Fig. 6. We observe that the activation probability decreases with increasing $K$. Note, however, that in the ODEs (10), the dependence on the label noise variance $\sigma^{2}$ disappears at leading order in $\eta$. This motivates us to compute the first-order correction in $\eta$ to determine how the optimal dropout rate depends on $\sigma^{2}$.
  • Unspecialized phase (introducing small-$\eta$ corrections): To investigate the role of label noise in determining the optimal activation rate, we derive perturbative corrections to the fixed point to leading order in $\eta$. In particular, we show analytically that the optimal activation probability takes the form
@@
    where $a(K,M)$ and $b(K,M)$ are coefficients we compute explicitly (see Fig. 7). Notably, we find that $b(K,M)<0$ for all values of $K$ and $M$ considered, confirming that the optimal activation probability decreases with increasing label noise.
  • Specialization transition: As anticipated, in the absence of dropout ($r=1$), the unspecialized phase is known to be linearly unstable [18], leading to specialization at long times—i.e., each student node aligns preferentially with one of the teacher nodes. Here, we investigate how dropout modifies this specialization transition. By performing a linear stability analysis around the plateau fixed point, we derive the phase diagram shown in Fig. 8, valid for $K=M$ and small $\eta$. Interestingly, we find that specialization occurs only for sufficiently large values of $r$ (i.e., low dropout). Below a critical threshold in $r$, the plateau phase becomes stable and the system does not transition to specialization. Finally, for $K=M$ and small $\eta$, we find that the optimal activation probability is given by
    $r^{*}\approx\min\left[1;\,1+\eta(a(M)+\sigma^{2}b(M))\right]\,$, (16)
    where $a(M)$ and $b(M)$ can be computed explicitly (see Fig. 9). We find $b(M)<0$ for all values of $M$ considered, confirming that dropout helps to mitigate label noise even in the specialized phase.
@@ 3 Results
In this section, we begin by discussing general insights obtained from the numerical integration of the ODEs (10). We then proceed with an analytical study of these equations, deriving explicit results for the optimal activation rate and the corresponding generalization performance across the different phases of the learning dynamics.
@@ 3.1 Numerical analysis of the dropout dynamics
We can therefore identify two competing objectives: 1) learning the signal $w_{1}^{*}$ effectively (favored by a low dropout rate) and 2) avoiding fitting the noise by promoting independence between $\tilde{w}_{1}$ and $\tilde{w}_{2}$ (favored by a higher dropout rate).
@@
Interestingly, dropout reduces the alignment of the weights with the signal $w_{1}^{*}$ (right panel in Fig. 2). This is expected, as dropout effectively reduces the number of training updates of each individual node. Despite the weaker signal representation, the benefit of increased independence between nodes outweighs this drawback, leading to superior generalization.
A key parameter in our setting is the variance $\sigma^{2}$ of the label noise. When training without dropout, the network is expected to be more prone to fitting this noise, which can degrade performance. In practical applications, dropout is commonly observed to mitigate noise-fitting and promote a more robust representation of the data.
Figure 3: Optimal activation rate $r^{*}$ as a function of the variance $\sigma^{2}$ of the label noise. Parameters: $K=2$, $M=1$, $\eta=1$, $\eta_{v}=0$, $\alpha=0.1$. The teacher weight vector $w^{*}$ is taken to be of unit norm, the student weights $w$ are initialized to zero. The second layer weights are fixed to one. The continuous line represents the optimal rate obtained by numerical optimization by using the full equations of motion. The crosses correspond to the analytical result in Eq. (24), obtained from a short-time expansion.
First, in Fig. 3 we examine the optimal dropout rate as a function of the label noise variance $\sigma^{2}$. Here, the crosses are obtained by numerical optimization of the full dynamical equations, while the continuous line corresponds to the asymptotic result derived analytically in Sec. 3.2. In the absence of label noise ($\sigma=0$), the optimal activation rate $r^{*}$ remains close to unity, indicating minimal dropout. Conversely, as the label noise increases, the optimal strategy shifts toward a higher dropout probability, corresponding to a lower activation probability.
As we will show in the following sections, the fact that the optimal activation rate $r^{*}$ decreases with increasing label noise holds robustly across the different phases of the learning dynamics.